Grade 12th passPhysical Chemistry

the energy of secound bohr orbital of the hydrogen atom is -328KJ/mol. hence the energy of fourth bhor orbit should be?

aayushi raghuvanshi
10 Years agoGrade 12th pass

1 Answer

Bhavya
10 Years ago
Dear student
En = E1 (Z2/n2)
E2 = E1 (12/22) = -328 (z=1,n=2)
E1 = -328*4
Now,
E4 = E1 (Z2/n2) {E1 = -328*4, n = 4, z = 1)
= -328*4 (12/42)
= -82 KJ/mol